OpenStudy (anonymous):

The sum of three numbers is 19 13/30. Two of the numbers are 5 2/3 and 6 1/6, What is the third number?

OpenStudy (raden):

let the third number is x x + 5 2/3 + 6 1/6 = 19 13/30 solve for x

OpenStudy (nbester98):

Another way to go but a bit more complex...turn this into a decimal x + 5.67 + 6.16 = 19.43 (div 13/30 to get decimal) x = 19.43-5.67-6.16 = 7.6 or x = 7 3/5